# Quantum to Classical Neural Network Transfer Learning Applied to Drug Toxicity Prediction

### v1

@@ 2.2 The Hadamard Test
The Hadamard test, shown in Figure 2, is another cornerstone quantum algorithm that allows one to determine $\text{Re}\braket{\psi}{U}{\psi}$ with error $\epsilon$, in $\mathcal{O}(\frac{1}{\epsilon^{2}})$ repetitions. However, by choosing the correct $U$, we can use this algorithm to produce the inner product between two vectors of choice, $\text{Re}\braket{\psi}{\phi}$. Since this gives a direction with the inner product unlike the swap test, we may dispense with the need for QPE as well as the need for separate quantum registers for the input data and parametric “weights”. Equations 1-4 show how $U$ can be chosen, for input data, |$\ket{\phi}$ and “weights”, |$\ket{\psi}$.
@@ 3.1 Hybrid Architecture
CNNs have been found to be successful for use in toxicity prediction tasks in the literature46, 47, 48, and are very popular QML analogs of classical architectures 49, 50, 51, 52, 53. The popularity of QCNNs partly stem from only needing to operate on small sections of input data at a time, rendering them feasible for NISQ devices. We employ a quantum-classical architecture where the first convolutional layer is replaced by a quantum circuit and is subsequently followed by classical pooling, a classical convolutional layer, and fully connected layers as depicted in Figure 4.
@@ 3.2 Tox21 Dataset
The Tox21 10k dataset 43, 44 is frequently used as a benchmark for evaluating ML models on their binary predictive ability for toxicity. This dataset comprises approximately 10,000 molecules represented as SMILES strings, each tagged with binary labels according to their response to 12 distinct assays. The Tox21 dataset suffers from significant imbalance, and although techniques such as SMOTE 54 can enhance performance, the primary objective of this study is to showcase the synergy between quantum and classical ML frameworks, rather than attaining the highest level of predictive accuracy. We specifically choose to create a sparse, binary feature map from each SMILES string in order to reduce overhead on the quantum component of the model. The computational benefit of binary features in a CNN is visually depicted in Figure 5. For each filter comprised of $w$ weights, there can exist a maximum of $2^{w}$ unique convolutions for binary data. Thus, for a single 2 $\times$ 2 filter there are only 16 unique dot-products that are required to fully stride across input data of any size. Additionally, the sparsity condition allows for the use of state-preparation techniques that do not scale exponentially 55, 56, which is a necessity along the road to achieving implementable quadratic complexity for matrix multiplication.
@@
Figure 5: Binary data in a CNN. The red and green sections of this example input data will result in an identical output. For our input data of size 400 $\times$ 57 and a single 2 $\times$ 2 filter, we only need to execute 15 circuits (the all-zero vector need not be computed), rather than 22,344 for each molecule.
In this study, we utilize the Binary-Encoded-SMILES (BES) approach 46, whereby each character within the SMILES string is depicted as a bit vector. 27 bits are used to encode the SMILES symbols and alphabet, and 30 bits are used to encode atomic properties. The vector of each character is joined with all other character vectors forming a grid, and is padded to the longest string in the dataset. The descriptors used to create this length 57 vector for each SMILES character are detailed in Table 1.
@@ 3.3 Quantum Transfer Learning
Figure 6: Quantum to classical neural network transfer pipeline. The yellow classical convolutional block L2-normalizes the input data covered by each filter to ensure a mathematically exact learning transition between quantum and classical computations.
@@ 4.1 Hardware
All scripting was done in Python 3.9.18. The quantum simulation package Pennylane 57 version 0.35.0 was used, as well as PyTorch 58 version 2.1.0 with CUDA Tookit 59 11.8 . An Intel i7-13700KF CPU, 12GB Nvidia GeForce RTX 3080Ti GPU, and 64GB of 3600MHz CL18 RAM were used for all computations.
@@ 4.3 Machine Learning
Binary cross entropy was used for the loss function and weights were updated with the Adam optimizer 60. A learning rate of 0.001 was used. Quantum weights were initialized from a uniform distribution between 00 and $2\pi$. To ensure fair comparisons between the quantum and classical models, all other classical parameters were randomly initialized and copied to be identical between the models. The exponential decay rate for the first and second moment estimations were set to 0.9 and 0.999, respectively for all training. Epsilon was set to 1.0 $\times$ $10^{-8}$. The bias component of the quantum convolutional layer was trained classically. The output of each hidden layer in the model architecture was activated with ReLU. Given the imbalanced nature of the binary dataset, we find it appropriate to measure performance with the area under the curve of the receiver operating characteristic (ROC-AUC) 61. This metric calculates the area under the curve of the plot of the true positive rate (Equation 6) versus the false positive rate (Equation 7). An ROC-AUC score of $1.0$ indicates the model predicts all samples correctly, while a score of $0.5$ indicates the model is no better than random chance.
@@ 4.4 Quantum Computing
Quantum states were prepared with the algorithm presented by Mottonen et al. 62 available in Pennylane, however in practice with this feature map, the algorithms by Veras et al. 55 and Gleinig et al. 56 would be better suited for use with NISQ devices. The variational component of the quantum circuit used three layers from Pennylane’s StronglyEntanglingLayers, inspired by Schuld et al. 63. The unitary matrices comprising the general rotation gates are show in Equations 8 and 9.
@@ 5 Results and Discussion
We compare our results to a fully classical model with identical architecture, except where the quantum convolutional filter is replaced by a classical filter of the same dimensions. It is shown in Table 2 that the QNN and CNN perform similarly for each assay. On the overall dataset, the quantum and classical models achieved an ROC-AUC of 0.692 and 0.689, respectively. These results demonstrate that the same accuracy as the classical $\mathcal{O}(n^{3})$ CNN can be achieved by the more efficient $\mathcal{O}(n^{2})$ quantum component of the model.
@@ 6 Conclusions
In summary, we present the first quantum-to-classical transfer learning framework where weights initially learned from a quantum circuit can be fine-tuned classically. We find quantum advantage by reducing the complexity of calculating matrix products from $\mathcal{O}(n^{3})$ to $\mathcal{O}(n^{2})$ using a modified version of the Hadamard test, which uses half the qubits of the previously proposed swap test for a quantum neuron and without the need for quantum phase estimation. We apply these methods to toxicity prediction using the Tox21 dataset, and show that our frameworks perform equivalently to a fully classical model with approximately the same number of parameters. As quantum devices improve, we hope this framework will provide an implementable polynomial speedup for machine learning networks, and consequently more efficiently navigate the chemical space to find life-saving pharmaceuticals.
@@ 9 Acknowledgements
We acknowledge the support provided by the NSF CCI grant (Award Number 2124511) and the National Institute of Health (Grant 1T32GM149438-01).

### v3

@@ 2.2 The Hadamard Test
The Hadamard test, shown in Figure 2, is another cornerstone quantum algorithm 46 that has found use in calculating inner products for fields ranging from chemistry 47, to physics 48, to mathematics 49. The test allows one to determine $\text{Re}\braket{\psi}{U}{\psi}$ with error $\epsilon$, in $\mathcal{O}(\frac{1}{\epsilon^{2}})$ repetitions. However, by choosing the correct $U$, we can use this algorithm to produce the inner product between two vectors of choice, $\text{Re}\braket{\psi}{\phi}$. Since this gives a direction with the inner product unlike the swap test, we may dispense with the need for QPE as well as the need for separate quantum registers for the input data and parametric “weights”. Equations 1-4 show how $U$ can be chosen, for input data, |$\ket{\phi}$ and “weights”, |$\ket{\psi}$.
@@ 3.1 Hybrid Architecture
CNNs have been found to be successful for use in toxicity prediction tasks in the literature50, 51, 52, and are very popular QML analogs of classical architectures 53, 54, 55, 56, 57. The popularity of QCNNs partly stem from only needing to operate on small sections of input data at a time, rendering them feasible for NISQ devices. We employ a quantum-classical architecture where the first convolutional layer is replaced by a quantum circuit and is subsequently followed by classical pooling, a classical convolutional layer, and a fully connected layer as depicted in Figure 4.
@@ 3.2 Tox21 Dataset
The Tox21 10k dataset 43, 44 is frequently used as a benchmark for evaluating ML models on their binary predictive ability for toxicity. This dataset comprises approximately 10,000 molecules represented as SMILES strings, each tagged with binary labels according to their response to 12 distinct assays. The Tox21 dataset suffers from significant imbalance, and although techniques such as SMOTE 58 can enhance performance, the primary objective of this study is to showcase the synergy between quantum and classical ML frameworks, rather than attaining the highest level of predictive accuracy. We specifically choose to create a sparse, binary feature map from each SMILES string in order to reduce overhead on the quantum component of the model. The computational benefit of binary features in a CNN is visually depicted in Figure 5. For each filter comprised of $w$ weights, there can exist a maximum of $2^{w}$ unique convolutions for binary data. Thus, for a single 2 $\times$ 2 filter there are only 16 unique dot-products that are required to fully stride across input data of any size. Additionally, the sparsity condition allows for the use of state-preparation techniques that do not scale exponentially 59, 60, which is a necessity along the road to achieving implementable quadratic complexity for matrix multiplication.
@@
Figure 5: Binary data in a CNN. The red and green sections of this example input data will result in an identical output, and thus only one of the two dot products need to be computed. Since the number of unique dot products for binary data is $2^{w}$, where $w$, is the number of weights in a filter, for our input data of size 400 $\times$ 57 and a single 2 $\times$ 2 filter, there are only $2^{2\times 2}$ different filters that will produce a unique dot product. Therefore, since the number of unique dot products is not dependent on the size of the input data, we only need to execute 15 circuits (the all-zero vector need not be computed), rather than the traditional 22,344 for each molecule. We note that the removal of non-unique calculations is applicable to both classical and quantum CNNs, however as quantum circuits are inherently more expensive than classical computations in the NISQ era, this process of removing duplicates becomes particularly valuable for QCNNs.
In this study, we utilize the Binary-Encoded-SMILES (BES) approach 50, whereby each character within the SMILES string is depicted as a bit vector. 27 bits are used to encode the SMILES symbols and alphabet, and 30 bits are used to encode atomic properties. The vector of each character is joined with all other character vectors forming a grid, and is padded to the longest string in the dataset. The descriptors used to create this length 57 vector for each SMILES character are detailed in Table 1.
@@ 3.3 Quantum Transfer Learning
Figure 6: Quantum to classical neural network transfer pipeline. The convolutional weights are derived from the learned unitary matrix $U_{\phi}$ by operating this on the ground state as shown in Equation 2. The real part of the amplitudes of the evolved state |$\ket{\phi}$ are identical to the normalized weights of a classical convolutional filter. The yellow classical convolutional block inherits the derived weights from the quantum state amplitudes. In order to maintain a mathematically exact learning transition between quantum and classical computations, this classical convolution L2-normalizes the input data covered by each filter. SLP indicates a Single Layer Perceptron. The exact architecture details are presented in Figure 4.
@@ 4.1 Hardware
All scripting was done in Python 3.9.18. The quantum simulation package Pennylane 61 version 0.35.0 was used, as well as PyTorch 62 version 2.1.0 with CUDA Toolkit 63 11.8 . An Intel i7-13700KF CPU, 12GB Nvidia GeForce RTX 3080Ti GPU, and 64GB of 3600MHz CL18 RAM were used for all computations.
@@ 4.3 Machine Learning
Binary cross entropy was used for the loss function and weights were updated with the Adam optimizer 64. A learning rate of 0.001 was used. Quantum weights were initialized from a uniform distribution between 00 and $2\pi$. To ensure fair comparisons between the quantum and classical models, all other classical parameters were randomly initialized and copied to be identical between the models. The exponential decay rate for the first and second moment estimations were set to 0.9 and 0.999, respectively for all training. Epsilon was set to 1.0 $\times$ $10^{-8}$. The bias component of the quantum convolutional layer was trained classically. The output of each hidden layer in the model architecture was activated with ReLU. Given the imbalanced nature of the binary dataset, we find it appropriate to measure performance with the area under the curve of the receiver operating characteristic (ROC-AUC) 65. This metric calculates the area under the curve of the plot of the true positive rate (Equation 6) versus the false positive rate (Equation 7). An ROC-AUC score of $1.0$ indicates the model predicts all samples correctly, while a score of $0.5$ indicates the model is no better than random chance.
@@ 4.4 Quantum Computing
Quantum states were prepared with the algorithm presented by Mottonen et al. 66 available in Pennylane, however in practice with this feature map, the algorithms by Veras et al. 59 and Gleinig et al. 60 would be better suited for use with NISQ devices. The variational component of the quantum circuit used three layers from Pennylane’s StronglyEntanglingLayers, inspired by Schuld et al. 67. The unitary matrices comprising the general rotation gates are show in Equations 8 and 9.
@@ 5 Results and Discussion
We compare our results to a fully classical model with identical architecture, except where the quantum convolutional filter is replaced by a classical filter of the same dimensions. It is shown in Table 2 that the hybrid quantum-classical neural network (QNN) and fully classical neural network (CNN) perform similarly for each assay. On the overall dataset, the quantum and classical models achieved an ROC-AUC of 0.692 and 0.689, respectively. These results demonstrate that the same accuracy as the fully classical $\mathcal{O}(n^{3})$ neural network can be achieved by the more efficient QNN with the $\mathcal{O}(n^{2})$ quantum component.
@@ 6 Conclusions
In summary, we present the first quantum-to-classical transfer learning framework where weights initially learned from a quantum circuit can be fine-tuned classically. Assuming efficient state preparation, we show a theoretical quantum advantage by reducing the complexity of calculating matrix products from $\mathcal{O}(n^{3})$ to $\mathcal{O}(n^{2})$ using a modified version of the Hadamard test, which uses half the qubits of the previously proposed swap test for a quantum neuron and without the need for quantum phase estimation. We apply these methods to toxicity prediction using the Tox21 dataset, and show that our frameworks perform equivalently to a fully classical model with approximately the same number of parameters. As quantum devices improve and the assumption of efficient state preparation becomes fulfilled, we hope this framework will provide an implementable polynomial speedup for machine learning networks, and consequently more efficiently navigate the chemical space to find life-saving pharmaceuticals.
@@ 9 Acknowledgements
We acknowledge financial support from the National Institute of Health Biophysical Training Grant (1T32GM149438-01) [AS], as well as the financial support provided by the CCI Phase I: National Science Foundation Center for Quantum Dynamics on Modular Quantum Devices (CQD-MQD) under Award Number 2124511 [VSB]. Furthermore, we acknowledge seed funding from Yale University through the National Science Foundation Engines Development Award: Advancing Quantum Technologies (CT) under Award Number 2302908 and high-performance computer time from the National Energy Research Scientific Computing Center and from the Yale University Faculty of Arts and Sciences High Performance Computing Center.